log6^(x+5)=log3^(x-4)
solve for x.

Answer by RAY100(1637)

log6^(x+5) = log 3 ^ (x-4)
.
(x+5) log6 = (x-4) log 3,,, from property of logs logx^3 = 3 logx
.
(x+5) * log6/log3 = (x-4)
.
(x+5) * 1.631 = (x-4)
.
1.631 x + 8.155 = x -4
.
.631 x = -12.155
.
x= (-19.263),,,,,answer
.
checking
.
log 6 ^ (-19.263 +5) = log 3 ^ (-19.263 -4)
.
log 6 ^ (-14.263) = log 3^ (-23.263)
.
-11.1 = -11.1 ,,,,,ok